OpenStudy (anonymous):

Find (f - h)(-3) if f(x) = (x + 1)^2; D = {Reals}, and h(x) = x; D = {Integers}

OpenStudy (anonymous):

f-h=(x+1)^2-x now let x=-3 what will u get?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

yes